Solve the Taylor's expression
F(x,y)=e^xcosy. (π,0)
1
Expert's answer
2020-03-04T11:17:36-0500

We find the partial derivatives of "f(x,y)= e^x cos(y)"

The partial derivatives of with respect to 'x' are "f{_x}= e^x cos(y)"

"f{_{xx}}= e^xcos(y)"

"f{_{xxx}}=e^x cos(y)......"

"\therefore f({{\pi},{0}})=f{_{x}}({\pi},0)=f{_{xx}}({\pi},0)....= e^{\pi}cos(0) = e^{\pi}"

The partial derivatives of "f(x,y)"

with respect to y are "f_{{y}}=e^x(-sin(y))"

"f_{{yy}}=e^x(-cos(y))"

"f_{{yyy}}=e^x(sin(y))...."

"f_{{y}}(\pi,0)=0=f_{{yyy}}(\pi,0)=f_{{yyyyy}}(\pi,0)...."

And "f_{{yy}}=-1"

"f_{{yyyy}}=1" ....

The partial derivatives of "\frac{\partial 2f}{\partial x\partial y}= 0"

Substituting these values in Taylor's

series we get

"f(x,y) = e^{{\pi}}+(x-\pi)e^{{\pi}}(1)+(x-\pi)^2e^{{\pi}}+0+y^2(-e^{{\pi}})+...."